dy 04: Solve Bernoullis equation dx + 3xy = y2x32
Answer #1

2 3 -yox (Bernoullis equation) dy da + 3xy 0 Compairing with we have dy da + P(x) y = Q(x) yn n = 2 4-2 = v V y dy du y2 da-322 2 2 V e - 3 x 2 -I (38²2).e tc 9 32 2 -1 (272+2) + C. y = (3x²+ 2) + ce 9 , C= arbitrary constant (: y = // is required